A Complete Bibliography of Publications in the Journal of Computational Geometry

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

13 August 2020 Version 1.01

Title word cross-reference

1 [BD16]. 1.88 [ABB^{+}16]. 2 [BEKT16, CW19]. 3 [HI20, HSW19, NPR16]. B_{2} [BD16]. d [Tan11]. k [BS14, MR13, PSW14]. L_{1} [CIW16, Wan20]. R^{3} [SV15, Sch16]. O(\sqrt{n}) [DH16]. Θ_{6} [BDD20].

-angulation [PSW14]. -connected [NPR16]. -manifold [BEKT16]. -manifolds [CW19, HI20, HSW19]. -means [MR13]. -non-degenerate [BS14]. -representability [Tan11]. -spanner [ABB^{+}16]. -string [BD16]. -VPG [BD16].

1.5D [FHKS16].

4-regular [BR15].

ABA [KP19]. ABA-free [KP19]. abstract [KP19]. actually [Dev11]. Adjacency [BEL^{+}16]. Adjacency-preserving [BEL^{+}16]. admit [PSW14]. affine [GGLT13]. algebraic [BB17, BL15]. Algorithm [Wan20, CS12, CW15, KS11, TW16]. Algorithms [FK20, MM15, ABdB^{+}15, AKP^{+}17, AMRW11, BEGM20]. Aligned [KP20]. All-pairs [CS19a]. almost [BCD^{+}16]. among [DKK^{+}17]. analysis [MR13]. anarchy [GZ12]. angle [CS12]. angle-constrained [CS12]. angles [ADD^{+}18]. angulation [PSW14]. antithickness [DW18]. application [BDD^{+}13]. Applications [Cha18, ADF18, BS14, DKT17, FK12, JR16, dBOS13]. approach [KP19]. approaching [NPR16]. Approximability [BM16]. Approximate [CS19b, Dum11]. Approximating [AdBK17, CKLS12]. approximation [BVX15, SV15, Sch16]. arc [DEG^{+}18]. area [AdBK17, BKR12, CW19, dBCC^{+}19]. areas [Kle18]. arrangements [DM20, Ort15]. Array [AD18]. Array-based [AD18]. aspects [Kön14]. average [CKLS12, DLS17]. axis [SV15, Sch16].

ball [MM15]. balls [BBM^{+}18, Epp17]. barcodes [BL15]. based [AD18, BB17, DKT17, DC19]. bases [BCFN17]. better [KPT17]. between [CW19]. bichromatic [AMR15]. bipartite [CF18]. black [dBRS12]. black-box [dBRS12]. blockers [PW10]. blue [BGL19]. bound [GJM12]. boundary [HI20]. bounds [BN12, DM20]. box [dBRS12, dBCC^{+}19]. Bregman [MR13].

Canonical [ADF18]. capturing [AU16]. case [GTH14b, MR13]. cell [TW16]. centerlines [BN12]. centers [Wan18]. Central [vKLS17]. characterizations [AMR15]. characterizing [ES15]. Chebyshev [Cha18]. Cheeger [GS15]. Chief [CR10]. chord [NPR16]. circle [BR15, GTH14a, GTH14b]. circle-representations [BR15]. circles [GTH14a, GTH14b]. circular [KN19]. Classifying [BLN17]. Cluster [WW20]. clustering [MR13, OPW20]. coins [ABC^{+}18]. collectors [DGGT16]. collinear [DDF^{+}18]. colorful [CKMU13]. coloring [BBM^{+}18, KP19]. Colouring [FMW12]. Combinatorial [BEGM20]. compact [AD18, HI20]. Competitive [BFvRV17]. complete [ADGL16, BG19, CF18, Ort15]. complex [Ker13]. complexes [ADGL16, BDT14, OPW20, Tan11]. Complexity [BDD20, BEGM20, BBM^{+}18, BGL19, DGGT16, GJM12, KLPS17, dBHT10]. Computational [Kön14, Cha18]. computations [DHLS13]. Computing [CSZ10, FHH^{+}17, FHH^{+}18, WN10, BKM^{+}18, CS12, CW15]. conditions [ACD^{+}16]. configurations [CCI^{+}19]. confluent [EHL^{+}16]. conjecture [BR15, Gru14]. Connected [LM11, NPR16]. connectivity [Epp19, dBRS12]. Conquer [Wan20]. Consistent [GNR16]. Constant [AMRW11, BVX15, DVW19]. constant-factor [BVX15]. Constant-work-space [AMRW11]. constrained [CS12, KN19]. constraints [BFvRV17]. construction [BB17, Epp11]. contact [AdCC^{+}12]. containers [AdBK17]. containing [KS15]. continuous [FHKS16]. convergence [CFL^{+}15]. convex [AdBK17, ABB^{+}16, BCCS10, BCD^{+}16, DGGT16, DHPT14, KKT15, NPPZ13, TW16, dBRS12]. convexity [FHKS17]. core [DDF^{+}14]. Cost [WW20]. counterexample [Gru14]. Counting [Wet17]. Cover [AdCC^{+}12, BP16, HPL12]. Covering [AAD^{+}14, dBCC^{+}19, BBG15, BVX15, GTH14a, GTH14b]. coverings [KP15]. crossing [Wet17]. crossing-free [Wet17]. crossings [KN19]. cube [FP18]. currents [IKV16]. curvature [DVW19]. curves [BH16, CW19, Hav11, Hav17, Sch16]. cusped [TW16]. cycle [BCFN17]. cycles [CCL16]. cylinder [ADF18].

data [AD18]. Declutter [BDWW18]. decomposing [KP15]. decomposition [GZ12, IKV16, LM14, TW16]. degenerate [BS14]. Degree [KPT17, BBD^{+}17]. Delaunay [BDT14, BCCS10, Dev11, DH16, KP20, dBRS12]. Delaunay-graphs [KP20]. denoising [BDWW18]. Density [AU16, DC19]. depth [ABDF^{+}11]. determined [FMW12]. Determining [HI20]. detour [WN10]. diagrams [BB17, BKM^{+}18, BEGM20, BDT14, BDD^{+}13, BKR12, DC19, SdMPS19]. dimension [KRW19, Kön14, Tan11]. dimensional [Cha18, GS15, Hav17]. dimensions [DK15]. directed [CCL16]. discrete [BM16, FK20, GS15]. discretization [FHKS16]. disjoint [Epp17]. disk [CS19b, KS15, PP11]. disks [BVX15, BBM^{+}18, Epp17, GKP^{+}16, KP20]. distance [BM16, CS19b, FK20, Kön14, vKLW20]. distances [NPPZ13]. distinct [NPPZ13]. divergences [MR13]. Divide [Wan20]. Divide-and-Conquer [Wan20]. Domains [Wan20, Wan18]. Drawing [DDF^{+}18, Kle18, BB17, EHL^{+}16]. drawings [ADF18, CF18, DEG^{+}18, KP20, NPR16]. dual [Ker13].

edge [ADD^{+}18, BF20, FHH^{+}18]. edge-length [BF20]. Editors [CD15, CR10, FL17]. Editors-in-Chief [CR10]. embedded [BCFN17]. Embedding [Ker13, Epp11]. embeddings [dBOS13]. emergence [GZ12]. empty [DHPT14]. encodings [CCI^{+}19]. endings [Epp10]. enumerating [Wet17]. equal [GTH14a, GTH14b]. estimating [KRW19]. estimation [DC19]. Euclidean [Dum11, Sol15, TW16]. exact [BBG^{+}18]. Expected [BDD20, DC19]. exponential [GJM12]. extension [BDD^{+}13].

face [Kle18]. faces [KKT15]. factor [BVX15, BCCS10, BCD^{+}16, CKLS12]. farthest [BDD^{+}13]. farthest-point [BDD^{+}13]. fast [Dev11, Epp11, LM14]. Fat [dBOS13]. feed [BDD^{+}13]. feed-link [BDD^{+}13]. filling [Hav11]. Finding [CCL16, KS15]. Fine [BBM^{+}18]. Fine-grained [BBM^{+}18]. fixed [KS11, Tan11]. fixed-parameter [KS11]. Flat [ADD^{+}18, IKV16, ADF18, Epp19, IKV13]. flip [Epp10]. Flow [DHLS13]. foldability [ADH^{+}20]. folded [AAD^{+}14, BH16]. foldings [ADD^{+}18, Epp19]. Forcing [Ort15, ACD^{+}16]. Foreword [CD15, FL17]. Four [KBK^{+}20, KPT17]. fragmentation [Hav11]. framework [DKT17]. Fréchet [FK20, vKLW20, BM16]. free [BDWW18, KP19, Wet17]. Function [WW20].

Gauss [MRZ16]. generalized [KP20]. geodesic [Wan18]. geometric [ABdB^{+}15, AMRW11, BP16, CS19a, CKLS12, FK12, Gru14, HPL12, Wet17, WN10]. geometry [Cha18, DKT17]. geometry-based [DKT17]. get [Dev11]. Good [PP11]. grained [BBM^{+}18]. graph [BKvRV18, KN19, WN10]. Graphs [BDD20, KBK^{+}20, ADD^{+}18, ADF18, AdCC^{+}12, BBD^{+}15, BR15, BD16, BCFN17, BF20, BCCS10, CCL16, CF18, CS19a, CS19b, CW15, CKLS12, DDF^{+}18, DW18, Epp10, Epp19, ES15, FK12, KP20, Kle18, LM11, NPR16, PP11, Wet17]. grouping [BBvK^{+}15]. growth [CDM^{+}16]. guarantees [AD18]. Guarding [BG19, GKKV14, FHKS16]. Guest [CD15, FL17].

Hales [Gru14]. Half [BDD20]. Half-

[BDD20]. Happy [Epp10]. hard [ADH^{+}20]. Hausdorff [DDF^{+}14, Kön14, vKLW20]. Hierarchical [WW20, CDM^{+}16]. high [DK15]. Higher [GS15]. Hilbert [BH16, Hav17]. hitting [KP19]. homology [BCFN17]. homotopy [CW19]. hulls [DGGT16, MM15, dBRS12]. hyper [Ker13]. hyper-rectangular [Ker13]. Hyperbolic [BDT14]. hyperbolicity [HI20]. hypergraphs [AU16, KP19]. Hyperorthogonal [BH16]. Hyperplane [FHKS17, DK15].

II [GTH14b]. implies [CDM^{+}16]. imprecise [Dev11, DHLS13, LM14]. Improved [BKM^{+}18, WW20, BBD^{+}15]. Incidences [BS14]. increasing [NPR16]. increasing-chord [NPR16]. incremental [Epp11]. Indeed [KBK^{+}20]. indexing [Het15]. Induced [BL15]. inequalities [Dum13, GS15]. integral [IKV16]. interference [DKK^{+}17]. intersection [CS19a, FK12]. intersections [MM15]. isocontours [BBG^{+}18]. isolating [GKP^{+}16]. issue [FL17].

Jewett [Gru14].

kernel [DC19]. Kinetic [dBRS12]. knot [SdMPS19].

label [OPW20]. labeling [GNR16]. landscapes [CFL^{+}15]. large [ABDF^{+}11]. largest [KS15]. layouts [KN19]. Lebesgue [BBG15]. length [BF20, FHH^{+}18]. lengths [ADD^{+}18]. light [Sol15]. line [ADF18]. linear [KS15, Sch16, dBHT10]. link [BBG^{+}18, BDD^{+}13, KLPS17]. little [Hav11]. local [BFvRV17, GKKV14]. logarithmic [KS15]. logspace [BEKT16]. Lombardi [DEG^{+}18]. Lovász [BR15]. low [Cha18]. low-dimensional [Cha18]. lower [DM20, GJM12]. LP [Gär15]. LP-type [Gär15].

made [BDT14]. Making [CKMU13]. Manhattan [Epp11, KS11]. manifold [BEKT16, KRW19]. manifolds [CW19, HI20, HSW19]. many [DDF^{+}18, Hav17, dBCC^{+}19]. maps [GNR16, dBHT10]. matching [BEGM20]. matchings [AMR15, BL15]. Maximizing [Epp17]. maximum [DHPT14, KKT15, WN10]. maxmin [FHH^{+}18]. means [MR13]. Measuring [CW19]. mechanical [GTH14a]. medial [SV15, Sch16]. medial-axis [SV15, Sch16]. median [DLS17]. Metric [Dum13, Koz14].

Minimax [KRW19]. Minimizing [KN19]. Minimum [BCFN17, BKR12, DHPT14, AdBK17, BBG^{+}18, FHH^{+}17, KS11, KLPS17]. minimum-area [AdBK17]. minimum-link [BBG^{+}18, KLPS17]. Minkowski [KKT15]. model [dBRS12]. modelled [DVW19]. models [GTH14a]. moving [DKK^{+}17]. multi [BP16, BVX15]. multi-cover [BP16]. multi-covering [BVX15]. multidimensional [CSZ10].

Near [Sch16]. Near-linear [Sch16]. Necessary [KBK^{+}20]. Network [BDD^{+}13, KS11]. networks [BBG^{+}18]. non [BS14, CCL16]. non-trivial [CCL16]. nonsimple [FHH^{+}17]. norm [IKV13, IKV16]. normed [MM15]. NP [ADH^{+}20, ADGL16, BG19]. NP-complete [ADGL16, BG19]. NP-hard [ADH^{+}20]. number [DM20, KKT15, NPPZ13].

obstacles [CW15]. octants [KP15]. off [AKP^{+}17]. offs [BKM^{+}18]. onion [LM14]. onions [LM14]. optimal [CS12, FHKS16, vKLW20]. Optimally [Epp11]. oracles [CS19b]. ordering [ADF18]. orientable [HI20, KGL19]. origami [ACD^{+}16, Epp19]. Orthogonal [BG19, EM14]. overlap [CDM^{+}16].

packing [AdBK17, GHK^{+}17]. Pages [KBK^{+}20]. pair [GZ12]. pairs [CS19a]. papers [FL17]. parallel [AMR15]. parameter [BDWW18, KS11]. parameterized [BGL19]. Part [GTH14a, GTH14b]. Partial [GTH14a, GTH14b]. partitions [BLN17, DHPT14, Ker13, dBOS13]. Path [Wan20, AEK^{+}16, CIW16, HP16, KLPS17]. paths [BBG^{+}18, CS19a, CS19b, GJM12]. Pattern [CDM^{+}16]. perimeter [FHH^{+}17]. periodic [ADF18]. persistence [BL15, CSZ10, CFL^{+}15, DC19]. perturbation [DKT17]. Placing [ABC^{+}18]. Planar [DEG^{+}18, KBK^{+}20, BR15, BD16, BF20, DDF^{+}18, Epp11, GHK^{+}17, Kle18, NPR16]. plane [ADD^{+}18, ABB^{+}16, BBD^{+}17, CW15, CIW16, Epp11, KPT17, KP20, WN10]. planes [MM15]. Point [Wan20, BDD^{+}13, CCI^{+}19, CIW16, FMW12, FHKS17, KS15, LM11, PSW14]. Points [ABDF^{+}11, ABB^{+}16, BGL19, Dev11, GKP^{+}16, LM14, dBCC^{+}19]. poly [DEG^{+}18]. poly-arc [DEG^{+}18]. polychromatic [KP19]. polygon [AKP^{+}17, HP16, NPPZ13]. Polygonal [Wan20, CW15, Wan18, dBOS13]. polygons [AdBK17, DDF^{+}14, Dum13, FHH^{+}17]. polyhedra [BCD^{+}16, EM14]. Polyhedral [MRZ16]. polyline [vKLW20]. polynomials [Cha18]. polyomino [BKR12]. polytope [GLMP16]. polytopes [DHPT14, KKT15, MRZ16]. position [ABB^{+}16]. practical [AD18, BDT14]. preprocess [Dev11]. preprocessing [LM14]. prescribed [ADD^{+}18, Kle18]. preserving [BEL^{+}16]. probabilistic [FHKS17]. problem [BBG15, DDF^{+}14, FHKS16, KS11]. problems [AMRW11, DKK^{+}17, Gär15, HPL12, KLPS17]. Progressive [ABdB^{+}15]. projection [DLS17]. projections [Koz14]. projective [TW16]. propagation [OPW20]. pseudo [KP20]. pseudo-disks [KP20]. pseudocircles [Ort15]. pseudohalfplanes [KP19]. pseudoline [DM20]. PTAS [FHKS16]. Ptolemaic [Het15].

quadrant [ABDF^{+}11]. Qualitative [DKT17]. quality [PP11]. Quasi [AMR15, BP16]. Quasi-parallel [AMR15]. quasi-uniform [BP16]. Queries [Wan20, AEK^{+}16, CIW16]. query [Dev11, KS15]. quickest [AEK^{+}16].

radii [Epp17]. Ramsey [Dum11]. Random [DK15, DH16, GLMP16]. range [AU16]. rates [KRW19]. ratio [BF20]. ratios [BBD^{+}15]. realistic [dBHT10]. Realization [Epp19, PP11]. recognition [BEKT16]. Recognizing [ADGL16]. rectangular [AdBK17, Ker13]. rectilinear [LM11]. Recursive [Hav11]. Recursively [JR16]. red [BGL19]. regular [BR15, JR16]. regularization [GJM12]. related [DKK^{+}17]. removal [Gär15]. representability [Tan11]. representation [BD16]. representations [BR15]. resample [BDWW18]. revisited [HPL12]. Rigid [ACD^{+}16, ADH^{+}20]. rigidity [GGLT13]. road [BBG^{+}18]. rotating [GNR16]. Routing [BKvRV18, BDD20, BFvRV17]. runaway [CDM^{+}16].

Sampling [Gär15, BP16]. Scalable [BBG^{+}18]. scale [IKV13]. Schnyder [KGL19]. search [DK15, GKKV14]. segment [AEK^{+}16]. segments [AMR15]. selected [FL17]. self [NPR16]. self-approaching [NPR16]. sensors [DKK^{+}17]. separability [FHKS17]. separated [GZ12]. separation [BGL19]. set [BP16, FMW12, HPL12]. sets [ACD^{+}16, BS14, FHKS17, KP19, LM11, PSW14]. shallow [KP19, Sol15]. shapes [AAD^{+}14]. shelf [ABC^{+}18]. Shortest [AEK^{+}16, HP16, Wan20, CCL16, CS19a, CS19b, CIW16]. shrinkable [ADGL16]. sided [KN19]. Silhouette [GLMP16]. silhouettes [CFL^{+}15]. similarity [CW19]. simple [AKP^{+}17, BB17, BR15, DDF^{+}14, EM14]. Simpler [KPT17]. Simplices [DVW19]. Simplicial [IKV13, OPW20, Tan11]. simplification [vKLW20]. small [dBCC^{+}19]. small-area [dBCC^{+}19]. smooth [Sch16]. Smoothed [DGGT16, MR13, dBHT10]. SoCG2016 [FL17]. solutions [AD18, FHKS16]. space [AKP^{+}17, AMRW11, BKM^{+}18, HP16, Hav11]. space-filling [Hav11]. spaces [DVW19, Koz14]. span [Epp11]. spanner [ABB^{+}16]. Spanners [FK12, BBD^{+}17, CS12, GZ12, KPT17]. spanning [BBD^{+}15]. sparse [GZ12]. sparsification [OPW20]. spatial [BEL^{+}16]. special [FL17]. Spectral [OPW20]. sphere [BCD^{+}16]. stability [BL15, FP18]. Stable [BEGM20]. Stable-matching [BEGM20]. Steiner [Sol15]. Steinitz [EM14]. Stochastic [CFL^{+}15]. storage [KS15]. straight [ADF18]. straight-line [ADF18]. stretch [BCCS10, BCD^{+}16, CKLS12]. Strict [EHL^{+}16]. strictly [TW16]. string [BD16]. structure [BBvK^{+}15, KGL19]. structures [AD18, dBRS12]. subarrangements [Ort15]. subdivisions [JR16]. sublinear [HP16]. Subquadratic [CCI^{+}19, SV15, WN10]. sum [Epp17, KKT15]. sums [MRZ16]. surface [BCFN17, TW16]. surface-embedded [BCFN17]. surfaces [CCL16, KGL19]. symbolic [DKT17]. symmetry [MRZ16]. systems [CDM^{+}16].

Terrain [BG19, ES15, FHKS16]. terrains [DHLS13, GKKV14, dBHT10]. tessellations [FP18]. their [BS14, BDD^{+}13]. theorem [FP18, GHK^{+}17]. theorems [Dum11, EM14]. theoretical [AD18]. there [Hav17]. Thickness [DW18]. three [Hav17, KKT15]. three-dimensional [Hav17]. tight [Epp11]. tile [CDM^{+}16]. tilings [Hav11]. Time [AKP^{+}17, BKM^{+}18, Dev11, KS15, Sch16, WN10]. Time-space [AKP^{+}17, BKM^{+}18]. Topological [CF18]. torus [ADF18, HI20]. trade [AKP^{+}17, BKM^{+}18]. trade-off [AKP^{+}17]. trade-offs [BKM^{+}18]. trajectories [vKLS17]. Trajectory [BBvK^{+}15]. translation [FK20]. tree [GHK^{+}17, SdMPS19]. tree-width [SdMPS19]. treemaps [BEL^{+}16]. Trees [WW20, DK15, Sol15]. treewidth [HSW19]. triangles [CKMU13, FMW12]. triangulated [HSW19]. triangulating [AKP^{+}17]. triangulation [Dev11, DH16]. triangulations [AD18, FHH^{+}18, dBRS12]. trivial [CCL16]. Tverberg [BLN17]. Two [CIW16, Wan20, DKT17, KN19]. Two-Point [Wan20, CIW16]. two-sided [KN19]. type [Gär15, Gru14].

unavoidable [BLN17]. unbounded

[Kön14]. uniform [BP16]. Unions [LM14]. unique [AMR15]. unit [BBM^{+}18, CS19b, GKP^{+}16, PP11]. unit-disk [CS19b]. universal [BBG15]. Upper [BN12]. using [GKP^{+}16, HP16, vKLW20].

Venn [BB17, BKR12]. vertex [NPPZ13]. vertices [ACD^{+}16, BCD^{+}16, DDF^{+}18]. via [BP16, BBG^{+}18, CW19, GKKV14]. virtual [PP11]. Visibility [dBHT10, AEK^{+}16, BKvRV18, CW15, DH16, ES15, PW10]. visualization [BBG^{+}18, dBOS13]. Voronoi [BKM^{+}18, BEGM20, BDT14]. VPG [BD16].

walk [DH16]. Weighted

[BP16, HPL12, CS19b, DLS17]. Welcome [CR10]. well [BH16, GZ12]. well-folded [BH16]. well-separated [GZ12]. Which [PSW14]. whose [BCD^{+}16]. width [SdMPS19]. witnesses [DGGT16]. woods [KGL19]. work [AMRW11]. Worst [MR13, DH16]. Worst-case [MR13].

Yao [BBD^{+}15].

References

Aichholzer:2014:CFS

[AAD^{+}14] Oswin Aichholzer, Greg Aloupis, Erik D. Demaine, Martin L. Demaine, Sándor P. Fekete, Michael Hoffmann, Anna Lubiw, Jack Snoeyink, and Andrew Winslow.

Covering folded shapes. Journal of Computational Geometry, 5(1):150–167, ???? 2014. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/160.

Amani:2016:PSP

[ABB^{+}16] Mahdi Amani, Ahmad Biniaz, Prosenjit Bose, Jean-Lou De Carufel, Anil Maheshwari, and Michiel Smid. A plane 1.88spanner for points in convex position. Journal of Computational Geometry, 7(1):520–539, ???? 2016. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/276.

Alt:2018:PYC

[ABC^{+}18] Helmut Alt, Kevin Buchin, Steven Chaplick, Otfried Cheong, Philipp Kindermann, Christian Knauer, and Fabian Stehn. Placing your coins on a shelf. Journal of Computational Geometry, 9(1):312–327, ???? 2018. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/361.

Alewijnse:2015:PGA

[ABdB^{+}15] Sander P. A. A. Alewijnse, Timur M. Bagautdinov, Mark de Berg, Quirijn W. Bouts, Alex P. ten Brink, Kevin Buchin, and Michel A. Westenberg. Progressive geometric algorithms. Journal of Computational Geometry, 6(2):72–92, ???? 2015.

CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/193.

Apfelbaum:2011:PLQ

[ABDF^{+}11] Roel Apfelbaum, Itay Ben-Dan, Stefan Felsner, Tillmann Miltzow, Rom Pinchasi, Torsten Ueckerdt, and Ran Ziv. Points with large quadrant depth. Journal of Computational Geometry, 2(1):128–143, ???? 2011. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/53.

Abel:2016:ROV

[ACD^{+}16] Zachary Abel, Jason Cantarella, Erik D. Demaine, David Eppstein, Thomas C. Hull, Jason S. Ku, Robert J. Lang, and Tomohiro Tachi. Rigid origami vertices: conditions and forcing sets. Journal of Computational Geometry, 7(1):171–184, ???? 2016. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/251.

Aleardi:2018:ABC

[AD18] Luca Castelli Aleardi and Olivier Devillers. Array-based compact data structures for triangulations: practical solutions with theoretical guarantees. Journal of Computational Geometry, 9(1):247–289, ???? 2018. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.

carleton.ca/jocg/index.php/
jocg/article/view/332.

Alt:2017:AMA

[AdBK17] Helmut Alt, Mark de Berg, and Christian Knauer. Approximating minimum-area rectangular and convex containers for packing convex polygons. Journal of Computational Geometry, 8(1):1–10, ???? 2017. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/289.

Atienza:2012:CCG

[AdCC^{+}12] Nieves Atienza, Natalia de Castro, Carmen Cortés, M. Ángeles Garrido, Clara I. Grima, Gregorio Hernández, Alberto Márquez, Auxiliadora Moreno-González, Martin Nöllenburg, José Ramón Portillo, Pedro Reyes, Jesús Valenzuela, Maria Trinidad Villar, and Alexander Wolff. Cover contact graphs. Journal of Computational Geometry, 3(1): 102–131, ???? 2012. CO-DEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/66.

Abel:2018:FFP

[ADD^{+}18] Zachary Abel, Erik D. Demaine, Martin L. Demaine, David Eppstein, Anna Lubiw, and Ryuhei Uehara. Flat foldings of plane graphs with prescribed angles and edge lengths. Journal of Computational Geometry, 9(1):74–93, ???? 2018.

CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/191. [Pag-

ing previously given as 71–91].

Aleardi:2018:COG

[ADF18] Luca Castelli Aleardi, Olivier Devillers, and Éric Fusy. Canonical ordering for graphs on the cylinder, with applications to periodic straight-line drawings on the flat cylinder and torus. Journal of Computational Geometry, 9(1):391–429, ???? 2018. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/360.

Attali:2016:RSC

[ADGL16] Dominique Attali, Olivier Devillers, Marc Glisse, and Sylvain Lazard. Recognizing shrinkable complexes is NP-complete. Journal of Computational Geometry, 7(1):430–443, ???? 2016. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/275.

Akitaya:2020:RFN

[ADH^{+}20] Hugo Akitaya, Erik Demaine, Takashi Horiyama, Thomas Hull, Jason Ku, and Tomohiro Tachi. Rigid foldability is NP-hard. Journal of Computational Geometry, 11(1):93–124, ???? 2020. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/403.

Arkin:2016:SPS

[AEK^{+}16] Esther M. Arkin, Alon Efrat, Christian Knauer, Joseph S. B. Mitchell, Valentin Polishchuk, Günter Rote, Lena Schlipf, and Topi Talvitie. Shortest path to a segment and quickest visibility queries. Journal of Computational Geometry, 7(2):77–100, ???? 2016. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/264.

Aronov:2017:TST

[AKP^{+}17] Boris Aronov, Matias Korman, Simon Pratt, André van Renssen, and Marcel Roeloffzen. Timespace trade-off algorithms for triangulating a simple polygon. Journal of Computational Geometry, 8(1):105–124, ???? 2017. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/307.

Asinowski:2015:QPS

[AMR15] Andrei Asinowski, Tillmann Miltzow, and Günter Rote. Quasi-parallel segments and characterizations of unique bichromatic matchings. Journal of Computational Geometry, 6(1): 185–219, ???? 2015. CO-DEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/134.

Asano:2011:CWS

[AMRW11] Tetsuo Asano, Wolfgang Mulzer, Günter Rote, and Yajun Wang. Constant-work-space algorithms for geometric problems. Journal of Computational Geometry, 2(1):46–68, ???? 2011. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/30.

Axenovich:2016:DRC

[AU16] Maria Axenovich and Torsten Ueckerdt. Density of range capturing hypergraphs. Journal of Computational Geometry, 7(1):1–21, ???? 2016. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/222.

Bannier:2017:NDS

[BB17] Arnaud Bannier and Nicolas Bodin. A new drawing for simple Venn diagrams based on algebraic construction. Journal of Computational Geometry, 8(1):153–173, ???? 2017. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/271.

Barba:2015:NIS

[BBD^{+}15] Luis Barba, Prosenjit Bose, Mirela Damian, Rolf Fagerberg, Wah Loon Keng, Joseph O’Rourke, André van Renssen, Perouz Taslakian, Sander Verdonschot, and Ge Xia. New and improved spanning ratios for Yao graphs. Journal of Computational Geometry, 6(2): 19–53, ???? 2015. CO-DEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/190.

Biniaz:2017:TPS

[BBD^{+}17] Ahmad Biniaz, Prosenjit Bose, Jean-Lou De Carufel, Cyril Gavoille, Anil Maheshwari, and Michiel Smid. Towards plane spanners of degree 3. Journal of Computational Geometry, 8(1):11–31, ???? 2017. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/295.

Baez:2015:LUC

[BBG15] John C. Baez, Karine Bagdasaryan, and Philip Gibbs. The Lebesgue universal covering problem. Journal of Computational Geometry, 6(1): 288–299, ???? 2015. CO-DEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/198.

Baum:2018:SEV

[BBG^{+}18] Moritz Baum, Thomas Bläsius, Andreas Gemsa, Ignaz Rutter, and Franziska Wegner. Scalable exact visualization of isocontours in road networks via minimum-link paths. Journal of Computational Geometry, 9(1):27–73, ???? 2018.

CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/313. [Pag-

ing previously given as 24–70].

Biro:2018:FGC

[BBM^{+}18] Csaba Biró, Édouard Bonnet, Dániel Marx, Tillmann Miltzow, and Paweł Rza̧żewski. Finegrained complexity of coloring unit disks and balls. Journal of Computational Geometry, 9(2):47–80, ???? 2018. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/414.

Buchin:2015:TGS

[BBvK^{+}15] Kevin Buchin, Maike Buchin, Marc van Kreveld, Bettina Speckmann, and Frank Staals. Trajectory grouping structure. Journal of Computational Geometry, 6(1):75–98, ???? 2015. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/161.

Bose:2010:SFC

[BCCS10] Prosenjit Bose, Paz Carmi, Sébastien Collette, and Michiel Smid. On the stretch factor of convex Delaunay graphs. Journal of Computational Geometry, 1(1):41–56, ???? 2010. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/5.

Bose:2016:SFC

[BCD^{+}16] Prosenjit Bose, Paz Carmi, Mirela Damian, Jean-Lou De Caroufel, Darryl Hill, Anil Maheshwari, Yuyang Liu, and Michiel Smid. On the stretch factor of convex polyhedra whose vertices are (almost) on a sphere. Journal of Computational Geometry, 7(1):444–472, ???? 2016. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/229.

Borradaile:2017:MCH

[BCFN17] Glencora Borradaile, Erin Wolf Chambers, Kyle Fox, and Amir Nayyeri. Minimum cycle and homology bases of surfaceembedded graphs. Journal of Computational Geometry, 8 (2):58–79, ???? 2017. CO-DEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/325.

Biedl:2016:SVR

[BD16] Therese Biedl and Martin Derka.

1-string B_{2}-VPG representation of planar graphs. Journal of Computational Geometry, 7(2):191–215, ???? 2016. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/296.

Bose:2013:NFP

[BDD^{+}13] Prosenjit Bose, Kai Dannies, Jean-Lou De Carufel, Christoph Doell, Carsten Grimm, Anil Maheshwari, Stefan Schirra, and Michiel Smid. Network farthestpoint diagrams and their application to feed-link network extension. Journal of Computational Geometry, 4(1):182–211, ???? 2013. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/131.

Bose:2020:ECR

[BDD20] Prosenjit Bose, Jean-Lou De Carufel, and Olivier Devillers. Expected complexity of routing in Θ_{6} and half-Θ_{6} graphs. Journal of Computational Geometry, 11(1):212–234, ???? 2020. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/484.

Bogdanov:2014:HDC

[BDT14] Mikhail Bogdanov, Olivier Devillers, and Monique Teillaud. Hyperbolic Delaunay complexes and Voronoi diagrams made practical. Journal of Computational Geometry, 5(1):56–85, ???? 2014. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/141.

Buchet:2018:DRT

[BDWW18] Mickaël Buchet, Tamal K. Dey, Jiayuan Wang, and Yusu Wang. Declutter and resample: towards parameter free denoising. Journal of Computational Geometry, 9(2):21–46, ???? 2018.

CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/402.

Barequet:2020:SMV

[BEGM20] Gill Barequet, David Eppstein, Michael Goodrich, and Nil Mamano. Stable-matching Voronoi diagrams: Combinatorial complexity and algorithms. Journal of Computational Geometry, 11(1):26–59, ???? 2020. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/428.

Burton:2016:MRL

[BEKT16] Benjamin A. Burton, Murray Elder, Arkadius Kalka, and Stephan Tillmann. 2-manifold recognition is in logspace. Journal of Computational Geometry, 7(1):70–85, ???? 2016. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/212.

Buchin:2016:APS

[BEL^{+}16] Kevin Buchin, David Eppstein, Maarten Löffler, Martin Nöllenburg, and Rodrigo I. Silveira. Adjacency-preserving spatial treemaps. Journal of Computational Geometry, 7(1): 100–122, ???? 2016. CO-DEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/211.

Borrazzo:2020:PEL

[BF20] Manuel Borrazzo and Fabrizio Frati. On the planar edge-length ratio of planar graphs. Journal of Computational Geometry, 11(1):137–155, ???? 2020. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/470.

Bose:2017:CLR

[BFvRV17] Prosenjit Bose, Rolf Fagerberg, André van Renssen, and Sander Verdonschot. Competitive local routing with constraints. Journal of Computational Geometry, 8(1):125–152, ???? 2017. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/288.

Bonnet:2019:OTG

[BG19] Édouard Bonnet and Panos Giannopoulos. Orthogonal terrain guarding is NP-complete. Journal of Computational Geometry, 10(2):21–44, ???? 2019. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/453.

Bonnet:2019:PCR

[BGL19] Edouard Bonnet, Panos Giannopoulos, and Michael Lampis. On the parameterized complexity of red–blue points separation. Journal of Computational Geometry, 10(1):181–206, ???? 2019. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/377.

Bos:2016:HWF

[BH16] Arie Bos and Herman Haverkort.

Hyperorthogonal well-folded Hilbert curves. Journal of Computational Geometry, 7(2):145–190, ???? 2016. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/269.

Banyassady:2018:ITS

[BKM^{+}18] Bahareh Banyassady, Matias Korman, Wolfgang Mulzer, André van Renssen, Marcel Roeloffzen, Paul Seiferth, and Yannik Stein. Improved time-space trade-offs for computing Voronoi diagrams. Journal of Computational Geometry, 9(1):191–212, ???? 2018. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/345.

Bultena:2012:MAP

[BKR12] Bette Bultena, Matthew Klimesh, and Frank Ruskey. Minimum area polyomino Venn diagrams. Journal of Computational Geometry, 3(1):154–167, ???? 2012. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/91.

Bose:2018:RVG

[BKvRV18] Prosenjit Bose, Matias Korman, André van Renssen, and Sander Verdonschot. Routing on the visibility graph. Journal of Computational Geometry, 9(1):430–453, ???? 2018. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/373.

Bauer:2015:IMA

[BL15] Ulrich Bauer and Michael Lesnick. Induced matchings and the algebraic stability of persistence barcodes. Journal of Computational Geometry, 6(2):162–191, ???? 2015. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/205.

Bukh:2017:CUT

[BLN17] Boris Bukh, Po-Shen Loh, and Gabriel Nivasch. Classifying unavoidable Tverberg partitions. Journal of Computational Geometry, 8(1):174–205, ???? 2017. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/308.

Bringmann:2016:ADF

[BM16] Karl Bringmann and Wolfgang Mulzer. Approximability of the discrete Fréchet distance. Journal of Computational Geometry, 7(2):46–76, ???? 2016. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/261.

Bukh:2012:UBC

[BN12] Boris Bukh and Gabriel Nivasch.

Upper bounds for centerlines. Journal of Computational Geometry, 3(1):20–30, ???? 2012. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/80.

Bansal:2016:WGS

[BP16] Nikhil Bansal and Kirk Pruhs.

Weighted geometric set multicover via quasi-uniform sampling. Journal of Computational Geometry, 7(1):221–236, ???? 2016. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/126.

Bekos:2015:CLC

[BR15] Michael A. Bekos and Chrysanthi N. Raftopoulou. On a conjecture of Lovász on circlerepresentations of simple 4regular planar graphs. Journal of Computational Geometry, 6(1):1–20, ???? 2015. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/120.

Basit:2014:IND

[BS14] Abdul Basit and Adam Sheffer.

Incidences with k-non-degenerate sets and their applications. Journal of Computational Geometry, 5(1):284–302, ???? 2014. CODEN ???? ISSN 1920-

180X. URL https://journals.

                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/138.

Bhowmick:2015:CFA

[BVX15] Santanu Bhowmick, Kasturi Varadarajan, and Shi-Ke Xue. A constant-factor approximation for multi-covering with disks. Journal of Computational Geometry, 6(1):220–234, ???? 2015. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/172.

Cardinal:2019:SEP

[CCI^{+}19] Jean Cardinal, Timothy M. Chan, John Iacono, Stefan Langerman, and Aurélien Ooms. Subquadratic encodings for point configurations. Journal of Computational Geometry, 10(2): 99–126, ???? 2019. CO-DEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/465.

Cabello:2016:FSN

[CCL16] Sergio Cabello, Éric Colin de Verdière, and Francis Lazarus. Finding shortest non-trivial cycles in directed graphs on surfaces. Journal of Computational Geometry, 7(1):123–148, ???? 2016. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/225.

Cheng:2015:GEF

[CD15] Siu-Wing Cheng and Olivier Devillers. Guest Editors’ foreword. Journal of Computational Geometry, 6(2):1, ???? 2015. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/206. Spe-

cial issue of selected papers from SoCG 2014, held in Kyoto, June 8–11, 2014.

Chen:2016:POI

[CDM^{+}16] Ho-Lin Chen, David Doty, Ján Maňuch, Arash Rafiey, and Ladislav Stacho. Pattern overlap implies runaway growth in hierarchical tile systems. Journal of Computational Geometry, 7(2):3–18, ???? 2016. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/255.

Cardinal:2018:TDC

[CF18] Jean Cardinal and Stefan Felsner. Topological drawings of complete bipartite graphs. Journal of Computational Geometry, 9(1):213–246, ???? 2018. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/347.

Chazal:2015:SCP

[CFL^{+}15] Frédéric Chazal, Brittany Terese Fasy, Fabrizio Lecci, Alessandro Rinaldo, and Larry Wasserman. Stochastic convergence of persistence landscapes and silhouettes. Journal of Computational Geometry, 6(2):140–161, ???? 2015. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/203.

Chan:2018:ACP

[Cha18] Timothy M. Chan. Applications of Chebyshev polynomials to lowdimensional computational geometry. Journal of Computational Geometry, 9(2):3–20, ???? 2018. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/391.

Chen:2016:TPL

[CIW16] Danny Z. Chen, Rajasekhar Inkulu, and Haitao Wang. Twopoint L_{1} shortest path queries in the plane. Journal of Computational Geometry, 7(1): 473–519, ???? 2016. CO-DEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/293.

Cheng:2012:AAS

[CKLS12] Siu-Wing Cheng, Christian Knauer, Stefan Langerman, and Michiel Smid. Approximating the average stretch factor of geometric graphs. Journal of Computational Geometry, 3(1):132–153, ???? 2012. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/87.

Cardinal:2013:MTC

[CKMU13] Jean Cardinal, Kolja Knauer, Piotr Micek, and Torsten Ueckerdt. Making triangles colorful. Journal of Computational Geometry, 4(1):240–246, ???? 2013. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/136.

Clarkson:2010:WEC

[CR10] Kenneth L. Clarkson and Günter Rote. Welcome from the Editorsin-Chief. Journal of Computational Geometry, 1(1):1–2, ???? 2010. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/20.

Carmi:2012:OAC

[CS12] Paz Carmi and Michiel Smid. An optimal algorithm for computing angle-constrained spanners. Journal of Computational Geometry, 3(1):196–221, ???? 2012. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/94.

Chan:2019:APS

[CS19a] Timothy M. Chan and Dimitrios Skrepetos. All-pairs shortest paths in geometric intersection graphs. Journal of Computational Geometry, 10 (1):27–41, ???? 2019. CO-DEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/384.

Chan:2019:ASP

[CS19b] Timothy M. Chan and Dimitrios Skrepetos. Approximate shortest paths and distance oracles in weighted unit-disk graphs. Journal of Computational Geometry, 10(2):3–20, ???? 2019. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/426.

Carlsson:2010:CMP

[CSZ10] Gunnar Carlsson, Gurjeet Singh, and Afra J. Zomorodian. Computing multidimensional persistence. Journal of Computational Geometry, 1(1):72–100, ???? 2010. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/19.

Chen:2015:NAC

[CW15] Danny Z. Chen and Haitao Wang. A new algorithm for computing visibility graphs of polygonal obstacles in the plane. Journal of Computational Geometry, 6(1):316–345, ???? 2015. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/232.

Chambers:2019:MSB

[CW19] Erin Wolf Chambers and Yusu Wang. Measuring similarity between curves on 2-manifolds via homotopy area. Journal of Computational Geometry, 10(1):96–126, ???? 2019.

CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/285.

deBerg:2019:CMP

[dBCC^{+}19] Mark de Berg, Sergio Cabello, Otfried Cheong, David Eppstein, and Christian Knauer. Covering many points with a small-area box. Journal of Computational Geometry, 10(1):207–222, ???? 2019. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/387.

deBerg:2010:VMR

[dBHT10] Mark de Berg, Herman Haverkort, and Constantinos P. Tsirogiannis. Visibility maps of realistic terrains have linear smoothed complexity. Journal of Computational Geometry, 1(1):57–71, ???? 2010. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/12.

deBerg:2013:FPP

[dBOS13] Mark de Berg, Krzysztof Onak, and Anastasios Sidiropoulos. Fat polygonal partitions with applications to visualization and embeddings. Journal of Computational Geometry, 4(1):212– 239, ???? 2013. CO-DEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/74.

deBerg:2012:KCH

[dBRS12] Mark de Berg, Marcel Roeloffzen, and Bettina Speckmann. Kinetic convex hulls, Delaunay triangulations and connectivity structures in the black-box model. Journal of Computational Geometry, 3(1):222–249, ???? 2012. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/89.

Divol:2019:DEP

[DC19] Vincent Divol and Frédéric Chazal. The density of expected persistence diagrams and its kernel based estimation. Journal of Computational Geometry, 10(2):127–153, ???? 2019. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/488.

Dorrigiv:2014:HCP

[DDF^{+}14] Reza Dorrigiv, Stephane Durocher, Arash Farzan, Robert Fraser, Alejandro López-Ortiz, J. Ian Munro, Alejandro Salinger, and Matthew Skala. The Hausdorff core problem on simple polygons. Journal of Computational Geometry, 5(1):14–40, ???? 2014. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/127.

DaLozzo:2018:DPG

[DDF^{+}18] Giordano Da Lozzo, Vida Dujmović, Fabrizio Frati, Tamara Mchedlidze, and Vincenzo Roselli. Drawing planar graphs with many collinear vertices. Journal of Computational Geometry, 9(1):94–130, ???? 2018. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/326.

Duncan:2018:PPA

[DEG^{+}18] Christian A. Duncan, David Eppstein, Michael T. Goodrich, Stephen G. Kobourov, Maarten Löffler, and Martin Nöllenburg. Planar and poly-arc Lombardi drawings. Journal of Computational Geometry, 9(1):328– 355, ???? 2018. CO-DEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/322.

Devillers:2011:DTI

[Dev11] Olivier Devillers. Delaunay triangulation of imprecise points: preprocess and actually get a fast query time. Journal of Computational Geometry, 2(1): 30–45, ???? 2011. CO-DEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/41.

Devillers:2016:SCC

[DGGT16] Olivier Devillers, Marc Glisse, Xavier Goaoc, and Rémy Thomasse. Smoothed complexity of convex hulls by witnesses and collectors. Journal of Computational Geometry, 7(2):101–144, ???? 2016.

CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/265.

Devillers:2016:WVW

[DH16] Olivier Devillers and Ross Hemsley. The worst visibility walk in a random Delaunay triangulation is O(\sqrt{n}). Journal of Computational Geometry, 7(1):332–359, ???? 2016. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/252.

Driemel:2013:FCI

[DHLS13] Anne Driemel, Herman Haverkort, Maarten Löffler, and Rodrigo I. Silveira. Flow computations on imprecise terrains. Journal of Computational Geometry, 4(1):38–78, ???? 2013. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/82.

Dumitrescu:2014:MCP

[DHPT14] Adrian Dumitrescu, Sariel Har-Peled, and Csaba D. Tóth. Minimum convex partitions and maximum empty polytopes. Journal of Computational Geometry, 5(1):86–103, ???? 2014. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/112.

Devroye:2015:RHS

[DK15] Luc Devroye and James King.

Random hyperplane search trees in high dimensions. Journal of Computational Geometry, 6(1):142–164, ???? 2015. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/67.

DeCarufel:2017:IAM

[DKK^{+}17] Jean-Lou De Carufel, Matthew J.

Katz, Matias Korman, André van Renssen, Marcel Roeloffzen, and Shakhar Smorodinsky. On interference among moving sensors and related problems. Journal of Computational Geometry, 8(1):32–46, ???? 2017. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/297.

Devillers:2017:QSP

[DKT17] Olivier Devillers, Menelaos I.

Karavelas, and Monique Teillaud. Qualitative symbolic perturbation: two applications of a new geometry-based perturbation framework. Journal of Computational Geometry, 8(1):282–315, ???? 2017. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/253.

Durocher:2017:PMW

[DLS17] Stephane Durocher, Alexandre Leblanc, and Matthew Skala.

The projection median as a weighted average. Journal of Computational Geometry, 8(1): 78–104, ???? 2017. CO-DEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/244.

Dumitrescu:2020:NLB

[DM20] Adrian Dumitrescu and Ritankar Mandal. New lower bounds for the number of pseudoline arrangements. Journal of Computational Geometry, 11(1):60–92, ???? 2020. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/435.

Dumitrescu:2011:AER

[Dum11] Adrian Dumitrescu. Approximate Euclidean Ramsey theorems. Journal of Computational Geometry, 2(1):16–29, ???? 2011. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/38.

Dumitrescu:2013:MIP

[Dum13] Adrian Dumitrescu. Metric inequalities for polygons. Journal of Computational Geometry, 4(1):79–93, ???? 2013. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/45.

Dyer:2019:SMS

[DVW19] Ramsay Dyer, Gert Vegter, and Mathijs Hubertus Maria Johannes Wintraecken. Simplices modelled on spaces of constant curvature. Journal of Computational Geometry, 10(1): 223–256, ???? 2019. CO-DEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/324.

Dujmovic:2018:TAG

[DW18] Vida Dujmovič and David R.

Wood. Thickness and antithickness of graphs. Journal of Computational Geometry, 9(1):356–386, ???? 2018. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/348.

Eppstein:2016:SCD

[EHL^{+}16] David Eppstein, Danny Holten, Maarten Löffler, Martin Nöllenburg, Bettina Speckmann, and Kevin Verbeek. Strict confluent drawing. Journal of Computational Geometry, 7(1):22–46, ???? 2016. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/174.

Eppstein:2014:STS

[EM14] David Eppstein and Elena Mumford. Steinitz theorems for simple orthogonal polyhedra. Journal of Computational Geometry, 5(1):179–244, ???? 2014.

CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/129.

Eppstein:2010:HEF

[Epp10] David Eppstein. Happy endings for flip graphs. Journal of Computational Geometry, 1(1):3–28, ???? 2010. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/21.

Eppstein:2011:OFI

[Epp11] David Eppstein. Optimally fast incremental Manhattan plane embedding and planar tight span construction. Journal of Computational Geometry, 2(1): 144–182, ???? 2011. CO-DEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/61.

Eppstein:2017:MSR

[Epp17] David Eppstein. Maximizing the sum of radii of disjoint balls or disks. Journal of Computational Geometry, 8(1):316–339, ???? 2017. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/286.

Eppstein:2019:RCG

[Epp19] David Eppstein. Realization and connectivity of the graphs of origami flat foldings. Journal of Computational Geometry, 10(1):257–280, ???? 2019.

CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/445.

Evans:2015:CTV

[ES15] William Evans and Noushin Saeedi. On characterizing terrain visibility graphs. Journal of Computational Geometry, 6(1):108–141, ???? 2015. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/130.

Fekete:2017:CNP

[FHH^{+}17] Sándor P. Fekete, Andreas Haas, Michael Hemmer, Michael Hoffmann, Irina Kostitsyna, Dominik Krupke, Florian Maurer, Joseph S. B. Mitchell, Arne Schmidt, Christiane Schmidt, and Julian Troegel. Computing nonsimple polygons of minimum perimeter. Journal of Computational Geometry, 8(1):340–365, ???? 2017. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/303.

Fekete:2018:CME

[FHH^{+}18] Sándor P. Fekete, Winfried Hellmann, Michael Hemmer, Arne Schmidt, and Julian Troegel. Computing maxmin edge length triangulations. Journal of Computational Geometry, 9(1): 1–26, ???? 2018. CO-DEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/

                                        jocg/article/view/319. [Pag-

ing previously given as 1–23].

Friedrichs:2016:CTG

[FHKS16] Stephan Friedrichs, Michael Hemmer, James King, and Christiane Schmidt. The continuous 1.5D terrain guarding problem: discretization, optimal solutions, and PTAS. Journal of Computational Geometry, 7(1):256–284, ???? 2016. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/242.

Fink:2017:HSC

[FHKS17] Martin Fink, John Hershberger, Nirman Kumar, and Subhash Suri. Hyperplane separability and convexity of probabilistic point sets. Journal of Computational Geometry, 8 (2):32–57, ???? 2017. CO-DEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/321.

Furer:2012:SGI

[FK12] Martin Fürer and Shiva Prasad Kasiviswanathan. Spanners for geometric intersection graphs with applications. Journal of Computational Geometry, 3(1): 31–64, ???? 2012. CO-DEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/31.

Filtser:2020:ADF

[FK20] Omrit Filtser and Matthew Katz. Algorithms for the discrete Fréchet distance under translation. Journal of Computational Geometry, 11(1):156–175, ???? 2020. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/409.

Fekete:2017:GEF

[FL17] Sándor P. Fekete and Anna Lubiw. Guest Editors’ foreword [special issue of selected papers from SoCG2016]. Journal of Computational Geometry, 8(2):1, ???? 2017. CO-DEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/368. Held

in Boston, MA, June 14–17, 2016.

Fabila-Monroy:2012:CTD

[FMW12] Ruy Fabila-Monroy and David R.

Wood. Colouring the triangles determined by a point set. Journal of Computational Geometry, 3(1):86–101, ???? 2012. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/78.

Frankl:2018:STC

[FP18] Peter Frankl and János Pach. A stability theorem on cube tessellations. Journal of Computational Geometry, 9(1):387– 390, ???? 2018. CO-DEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/398.

Gartner:2015:SRL

[Gär15] Bernd Gärtner. Sampling with removal in LP-type problems. Journal of Computational Geometry, 6(2):93–112, ???? 2015. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/197.

Gortler:2013:AR

[GGLT13] Steven J. Gortler, Craig Gotsman, Ligang Liu, and Dylan P. Thurston. On affine rigidity. Journal of Computational Geometry, 4(1):160–181, ???? 2013. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/49.

Geyer:2017:PTP

[GHK^{+}17] Markus Geyer, Michael Hoffmann, Michael Kaufmann, Vincent Kusters, and Csaba D. Tóth. The planar tree packing theorem. Journal of Computational Geometry, 8(2):109–177, ???? 2017. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/341.

Gartner:2012:ELB

[GJM12] Bernd Gärtner, Martin Jaggi, and Clément Maria. An exponential lower bound on the complexity of regularization paths.

Journal of Computational Geometry, 3(1):168–195, ???? 2012. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/88.

Gibson:2014:GTL

[GKKV14] Matt Gibson, Gaurav Kanade, Erik Krohn, and Kasturi Varadarajan. Guarding terrains via local search. Journal of Computational Geometry, 5(1):168–178, ???? 2014. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/128.

Gibson:2016:IPU

[GKP^{+}16] Matt Gibson, Gaurav Kanade, Rainer Penninger, Kasturi Varadarajan, and Ivo Vigan. On isolating points using unit disks. Journal of Computational Geometry, 7(1):540–557, ???? 2016. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/237.

Glisse:2016:SRP

[GLMP16] Marc Glisse, Sylvain Lazard, Julien Michel, and Marc Pouget. Silhouette of a random polytope. Journal of Computational Geometry, 7(1):86–99, ???? 2016. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/162.

Gemsa:2016:CLR

[GNR16] Andreas Gemsa, Martin Nöllenburg, and Ignaz Rutter. Consistent labeling of rotating maps. Journal of Computational Geometry, 7(1):308–331, ???? 2016. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/184.

Gruslys:2014:CGH

[Gru14] Vytautas Gruslys. A counterexample to a geometric Hales– Jewett type conjecture. Journal of Computational Geometry, 5(1):245–249, ???? 2014. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/176.

Gundert:2015:HDD

[GS15] Anna Gundert and May Szedlák. Higher dimensional discrete Cheeger inequalities. Journal of Computational Geometry, 6(2):54–71, ???? 2015. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/192.

Gaspar:2014:PCCa

[GTH14a] Zsolt Gáspár, Tibor Tarnai, and Krisztián Hincz. Partial covering of a circle by equal circles. Part I: the mechanical models. Journal of Computational Geometry, 5(1):104–125, ???? 2014. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.

carleton.ca/jocg/index.php/
jocg/article/view/108.

Gaspar:2014:PCCb

[GTH14b] Zsolt Gáspár, Tibor Tarnai, and Krisztián Hincz. Partial covering of a circle by equal circles. Part II: the case of 5 circles. Journal of Computational Geometry, 5(1):126–149, ???? 2014. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/109.

Gao:2012:ESS

[GZ12] Jie Gao and Dengpan Zhou. The emergence of sparse spanners and well-separated pair decomposition under anarchy. Journal of Computational Geometry, 3(1):1–19, ???? 2012. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/63.

Haverkort:2011:RTS

[Hav11] Herman Haverkort. Recursive tilings and space-filling curves with little fragmentation. Journal of Computational Geometry, 2(1):92–127, ???? 2011. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/68.

Haverkort:2017:HMT

[Hav17] Herman Haverkort. How many three-dimensional Hilbert curves are there? Journal of Computational Geometry, 8(1):

206–281, ???? 2017. CO-DEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/298.

Hetland:2015:PI

[Het15] Magnus Lie Hetland. Ptolemaic indexing. Journal of Computational Geometry, 6(1): 165–184, ???? 2015. CO-DEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/42.

Haraway:2020:DHC

[HI20] Robert Cyrus Haraway and III. Determining hyperbolicity of compact orientable 3-manifolds with torus boundary. Journal of Computational Geometry, 11(1):125–136, ???? 2020. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/433.

Har-Peled:2016:SPP

[HP16] Sariel Har-Peled. Shortest path in a polygon using sublinear space. Journal of Computational Geometry, 7(2):19–45, ???? 2016. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/256.

Har-Peled:2012:WGS

[HPL12] Sariel Har-Peled and Mira Lee. Weighted geometric set cover problems revisited. Journal of Computational Geometry, 3 (1):65–85, ???? 2012. CO-DEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/77.

Huszar:2019:TTM

[HSW19] Kristóf Huszár, Jonathan Spreer, and Uli Wagner. On the treewidth of triangulated 3manifolds. Journal of Computational Geometry, 10(2): 70–98, ???? 2019. CO-DEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/479.

Ibrahim:2013:SFN

[IKV13] Sharif Ibrahim, Bala Krishnamoorthy, and Kevin R. Vixie. Simplicial flat norm with scale. Journal of Computational Geometry, 4(1):133–159, ???? 2013. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/85.

Ibrahim:2016:FND

[IKV16] Sharif Ibrahim, Bala Krishnamoorthy, and Kevin R. Vixie. Flat norm decomposition of integral currents. Journal of Computational Geometry, 7(1): 285–307, ???? 2016. CO-DEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/199.

Jaume:2016:RRS

[JR16] Rafel Jaume and Günter Rote. Recursively regular subdivisions and applications. Journal of Computational Geometry, 7(1): 185–220, ???? 2016. CO-DEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/186.

Kaufmann:2020:FPI

[KBK^{+}20] Michael Kaufmann, Michael Bekos, Fabian Klute, Sergey Pupyrev, Chrysanthi Raftopoulou, and Torsten Ueckerdt. Four pages are indeed necessary for planar graphs. Journal of Computational Geometry, 11(1): 332–353, ???? 2020. CO-DEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/504.

Kerber:2013:EDC

[Ker13] Michael Kerber. Embedding the dual complex of hyperrectangular partitions. Journal of Computational Geometry, 4(1):13–37, ???? 2013. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/104.

Knauer:2019:SSW

[KGL19] Kolja Knauer, Daniel Gonçalves, and Benjamin Leveque. On the structure of Schnyder woods on orientable surfaces. Journal of Computational Geometry, 10(1):127–163, ???? 2019.

CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/386.

Karavelas:2015:MNF

[KKT15] Menelaos I. Karavelas, Christos Konaxis, and Eleni Tzanaki. The maximum number of faces of the Minkowski sum of three convex polytopes. Journal of Computational Geometry, 6(1): 21–74, ???? 2015. CO-DEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/142.

Kleist:2018:DPG

[Kle18] Linda Kleist. Drawing planar graphs with prescribed face areas. Journal of Computational Geometry, 9(1):290–311, ???? 2018. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/351.

Kostitsyna:2017:CML

[KLPS17] Irina Kostitsyna, Maarten Löffler, Valentin Polishchuk, and Frank Staals. On the complexity of minimum-link path problems. Journal of Computational Geometry, 8(2):80–108, ???? 2017. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/328.

Klute:2019:MCC

[KN19] Fabian Klute and Martin Nöllenburg.

Minimizing crossings in constrained two-sided circular graph layouts. Journal of Computational Geometry, 10(2): 45–69, ???? 2019. CO-DEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/476.

Konig:2014:CAH

[Kön14] Stefan König. Computational aspects of the Hausdorff distance in unbounded dimension. Journal of Computational Geometry, 5(1):250–274, ???? 2014. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/156.

Kozdoba:2014:PMS

[Koz14] Mark Kozdoba. On projections of metric spaces. Journal of Computational Geometry, 5(1):275–283, ???? 2014. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/72.

Keszegh:2015:MDC

[KP15] Balázs Keszegh and Dömötör Pálvölgyi. More on decomposing coverings by octants. Journal of Computational Geometry, 6(1):300–315, ???? 2015. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/224.

Keszegh:2019:AAP

[KP19] Balázs Keszegh and Dömötör Pálvölgyi. An abstract approach to polychromatic coloring: shallow hitting sets in ABA-free hypergraphs and pseudohalfplanes. Journal of Computational Geometry, 10(1):1–26, ???? 2019. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/349.

Keszegh:2020:APD

[KP20] Balázs Keszegh and Dömötör Pálvölgyi. Aligned plane drawings of the generalized Delaunaygraphs for pseudo-disks. Journal of Computational Geometry, 11(1):354–370, ???? 2020. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/475.

Kanj:2017:DFP

[KPT17] Iyad Kanj, Ljubomir Perković, and Duru Türkoğlu. Degree four plane spanners: Simpler and better. Journal of Computational Geometry, 8(2):3–31, ???? 2017. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/320.

Kim:2019:MRE

[KRW19] Jisu Kim, Alessandro Rinaldo, and Larry Wasserman. Minimax rates for estimating the dimension of a manifold. Journal of Computational Geometry, 10(1):42–95, ???? 2019.

CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/278.

Knauer:2011:FPA

[KS11] Christian Knauer and Andreas Spillner. A fixed-parameter algorithm for the minimum Manhattan network problem. Journal of Computational Geometry, 2(1):189–204, ???? 2011. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/43.

Kaminker:2015:FLD

[KS15] Tal Kaminker and Micha Sharir. Finding the largest disk containing a query point in logarithmic time with linear storage. Journal of Computational Geometry, 6(2):3–18, ???? 2015. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/188.

Loffler:2011:CRG

[LM11] Maarten Löffler and Elena Mumford. Connected rectilinear graphs on point sets. Journal of Computational Geometry, 2(1):1–15, ???? 2011. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/40.

Loffler:2014:UOP

[LM14] Maarten Löffler and Wolfgang Mulzer. Unions of onions: preprocessing imprecise points for fast onion decomposition. Journal of Computational Geometry, 5(1):1–13, ???? 2014. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/140.

Martin:2015:ABH

[MM15] Pedro Mart´ ın and Horst Martini. Algorithms for ball hulls and ball intersections in normed planes. Journal of Computational Geometry, 6(1):99–107, ???? 2015. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/187.

Manthey:2013:WCS

[MR13] Bodo Manthey and Heiko Röglin.

Worst-case and smoothed analysis of k-means clustering with Bregman divergences. Journal of Computational Geometry, 4(1):94–132, ???? 2013. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/39.

Malikiosis:2016:PGS

[MRZ16] Romanos-Diogenes Diogenes Malikiosis, Sinai Robins, and Yichi Zhang. Polyhedral Gauss sums, and polytopes with symmetry. Journal of Computational Geometry, 7(1):149–170, ???? 2016. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/231.

Nivasch:2013:NDD

[NPPZ13] Gabriel Nivasch, János Pach, Rom Pinchasi, and Shira Zerbib. The number of distinct distances from a vertex of a convex polygon. Journal of Computational Geometry, 4(1):1–12, ???? 2013. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/103.

Nollenburg:2016:SAI

[NPR16] Martin Nöllenburg, Roman Prutkin, and Ignaz Rutter. On self-approaching and increasingchord drawings of 3-connected planar graphs. Journal of Computational Geometry, 7(1): 47–69, ???? 2016. CO-DEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/223.

Osting:2020:SSS

[OPW20] Braxton Osting, Sourabh Palande, and Bei Wang. Spectral sparsification of simplicial complexes for clustering and label propagation. Journal of Computational Geometry, 11(1):176–211, ???? 2020. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/417.

Ortner:2015:FSC

[Ort15] Ronald Ortner. Forcing subarrangements in complete arrangements of pseudocircles. Journal of Computational Geometry, 6(1):235–248, ???? 2015. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/37.

Pemmaraju:2011:GQV

[PP11] Sriram V. Pemmaraju and Imran A. Pirwani. Good quality virtual realization of unit disk graphs. Journal of Computational Geometry, 2(1): 69–91, ???? 2011. CO-DEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/35.

Payne:2014:WPS

[PSW14] Michael S. Payne, Jens M.

Schmidt, and David R. Wood. Which point sets admit a kangulation? Journal of Computational Geometry, 5(1): 41–55, ???? 2014. CO-DEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/92.

Por:2010:VB

[PW10] Attila Pór and David R. Wood.

On visibility and blockers. Journal of Computational Geometry, 1(1):29–40, ???? 2010. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/24.

Scheffer:2016:NLT

[Sch16] Christian Scheffer. Near-linear time medial-axis approximation of smooth curves in R^{3}. Journal of Computational Geometry, 7(1):360–429, ???? 2016. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/238.

Schleimer:2019:TWK

[SdMPS19] Saul Schleimer, Arnaud de Mesmay, Jessica Purcell, and Eric Sedgwick. On the tree-width of knot diagrams. Journal of Computational Geometry, 10(1): 164–180, ???? 2019. CO-DEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/408.

Solomon:2015:ESS

[Sol15] Shay Solomon. Euclidean Steiner shallow–light trees. Journal of Computational Geometry, 6(2): 113–139, ???? 2015. CO-DEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/202.

Scheffer:2015:SMA

[SV15] Christian Scheffer and Jan Vahrenhold. Subquadratic medial-axis approximation in R^{3}. Journal of Computational Geometry, 6(1):249–287, ???? 2015. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/148.

Tancer:2011:RSC

[Tan11] Martin Tancer. d-representability of simplicial complexes of fixed dimension. Journal of Computational Geometry, 2(1):183– 188, ???? 2011. CO-DEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/73.

Tillmann:2016:AEC

[TW16] Stephan Tillmann and Sampson Wong. An algorithm for the Euclidean cell decomposition of a cusped strictly convex projective surface. Journal of Computational Geometry, 7(1):237–255, ???? 2016. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/241.

vanKreveld:2017:CT

[vKLS17] Marc van Kreveld, Maarten Löffler, and Frank Staals. Central trajectories. Journal of Computational Geometry, 8(1): 366–386, ???? 2017. CO-DEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/302.

vanKreveld:2020:OPS

[vKLW20] Marc van Kreveld, Maarten Löffler, and Lionov Wiratma. On optimal polyline simplification using the Hausdorff and Fréchet distance. Journal of Computational Geometry, 11(1):1–25, ???? 2020.

CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/415.

Wang:2018:GCP

[Wan18] Haitao Wang. On the geodesic centers of polygonal domains. Journal of Computational Geometry, 9(1):131–190, ???? 2018. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/290.

Wang:2020:DCA

[Wan20] Haitao Wang. A divide-and-conquer algorithm for two-point L_{1} shortest path queries in polygonal domains. Journal of Computational Geometry, 11(1): 235–282, ???? 2020. CO-DEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/462.

Wettstein:2017:CEC

[Wet17]  Manuel Wettstein.  Count-

ing and enumerating crossingfree geometric graphs. Journal of Computational Geometry, 8(1):47–77, ???? 2017. CODEN ???? ISSN 1920-

                                        180X. URL https://journals.
                                        carleton.ca/jocg/index.php/
                                        jocg/article/view/280.

Wulff-Nilsen:2010:CMD

[WN10] Christian Wulff-Nilsen. Computing the maximum detour of a plane geometric graph in subquadratic time. Journal of Computational Geometry, 1(1):101–122, ???? 2010. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/6.

Wang:2020:ICF

[WW20] Dingkang Wang and Yusu Wang.

An improved cost function for hierarchical cluster trees. Journal of Computational Geometry, 11(1):283–331, ???? 2020. CODEN ???? ISSN 1920-

180X. URL https://journals.
carleton.ca/jocg/index.php/
jocg/article/view/431.